an object is dropped from a 42 m tall building. how long does it take to reach the ground
1
Expert's answer
2018-11-01T16:00:08-0400

When object was dropped from some height, it displacement

h=v_i t+(at^2)/2=(gt^2)/2

So

t=√(2h/g)=√((2×42)/9.8)=2.9 s